In a basketball game carl made 16 shots. Each of his shots was worth either 2 or 3 points and jerry scored a total of 39 points. Let x represent the number of two point shots and y represents the number of 3 point shots. Write the system of equations in terms of x and y.
Answer by Boreal(15235)

x=2 pointers
y=3 pointers
x+y=16
2x+3y=39
========
2x+2y=-32
-2x-3y=-39
-y=-7
y=7 (21 points)
x=9 (18 points)